OpenStudy (anonymous):

What is the circumference of a circle with an area of 64(pi) in^2 ? a. 8(pi)in b. 16(pi)in c. 1024(pi)in d. 32(pi)in e. none of these

OpenStudy (shinalcantara):

can you recall the formula in finding area and circumference of a circle?

OpenStudy (cj49):

circumference= 2 pi r

OpenStudy (anonymous):

2 pi r

OpenStudy (anonymous):

can you show me the work leading to the answer?

OpenStudy (shinalcantara):

that's for circumference. for area, it's pi r squared. since you're given the area of the circle, substitute it to the formula for area. what you've got?

OpenStudy (cj49):

A is the ans Area= pi x r^2 64= pi x r^2 r=8 pi

OpenStudy (shinalcantara):

what's the value of your 'r'?

OpenStudy (princeharryyy):

nice work @cj49

OpenStudy (cj49):

lol thnx mate @princeharryyy

OpenStudy (princeharryyy):

:p

OpenStudy (shinalcantara):

A=(pi)*r^2 C=2(pi)r since you're given the area to be 64 pi, then using the area formula, you'll get the value of r as 8.. substitute the value to the formula in finding circumference. then you'll have 2(pi)*8. therefore your circumference is 16 pi.